# SN 2021lwz: Another Exotic Luminous and Fast Evolving Optical Type Ic Broad-Lined Supernova ?

### v1

@@ 5 Discussion
Based on our previous modeling and analysis it is not clear what is the nature of SN 2021lwz. This very fast evolving explosion (t days, rest-frame), with a low ejecta mass (), is drawn from a very low-brightness ( mag), low mass () dwarf galaxy of SFR () about ten times larger than the typical star-forming galaxies.
| Object | R.A. | Decl. | Type | z | trise,rf (a) | band | (b) | (c) | Ref. |
@@
To discuss SN 2021lwz in a broader context, in Figure 16 we show the rest-frame light curves of the different types of transients listed in Table 3; namely SN 2018bgv (SLSN-I), SN 1998bw (peculiar Ic-BL), AT 2020xnd (LFBOT), SN 2014ft (USSN), iPTF16asu (SL Ic-BL) and SN 2018gep (SL Ic-BL). As it is done for SN 2021lwz, if the photometry is not already corrected for the extinction contribution from our Galaxy, it is done using Schlafly & Finkbeiner (2011). As much as possible the filters were chosen by taking into account the redshift of each source so that the rest-frame wavelength ranges covered overlap strongly with each other.
@@ 5.1 H-poor SLSNe
SN 2018bgv is the fastest evolving ( days rest-frame) Type-W SLSN-I reported in the literature. It peaks close to mag. It is therefore very interesting for comparisons with SN 2021lwz. While we didn’t model it with our magnetar model, the light curve and a spectral sequence have been studied in detail by Lunnan et al. (2020). From black-body fits analysis of the light curve the authors report a photosphere temperature close to 11,000 K and a black-body radius close to 2 cm, a few days after peak, i.e. a photopshere with a radius about twice larger than, and a temperature about similar to, those of SN 2021lwz photosphere at similar phases. Close to peak the velocity of the photosphere of SN 2018bgv appears to be about 1.5 times the velocity estimated for SN 2021lwz. A fit of the light curve with the magnetar spin-down model using mosfit infers an ejecta of mass, , i.e. about 5 times higher than the one of SN 2021lwz as from our modeling estimate, hence suggesting the conditions behind the explosions of SN 2018bgv and SN 2021lwz are different.
As discussed in Section 3.1 SN 2018bgv provides a relatively good match with SN 2021lwz close to maximum light, but this is obtained at a lower redshift z than the redshift of SN 2021lwz. This is mainly because SNID tries to adjust the archetypal W-shape feature attributed to O ii, before and close to maximum light, observed in many Type I SLSNe. Indeed, and as discussed in Section 4.1, for H-poor SLSNe the typical absorption spectral features expected to be observed before maximum light, between 4000 and 5000 Å (see Könyves-Tóth & Seli 2023, Figure 1), in Type 15bn (e.g. Fe iii) and in Type W (e.g. O ii), are not detected in the pre-maximum spectra of SN 2021lwz. Instead, C ii, O i, Na i, Mg ii, Si ii, P iii, Ca ii, Ti ii and Fe ii are included to reproduce the shape of these spectra, and this despite the fact that from the estimates of the photosphere temperature K and the photosphere velocities kms before maximum light, one would expect features from a W-Type SLSN-I (see Könyves-Tóth 2022, Figure 7). We point out, however, that the conditions for these lines to appear are likely limited to a more stringent range of temperature K once spectral modeling effects, which are not included in the SYN++ modeling, are taken into account (see Saito et al. 2024).
@@ 5.2 LFBOTs
The LFBOT AT 2020xnd and SN 2021lwz are peaking at about similar absolute magnitudes (see Figure 16), but AT 2020xnd is a faster evolving event than SN 2021lwz, mainly because it has a higher initial pulsar luminosity, assuming a magnetar model. Two spectra of AT 2020xnd at phases close to +5 days and +7 days after maximum light are publicly available in the WISeREP repository 171717URL to AT 2020xnd public spectra in WISeREP, https://www.wiserep.org/object/18199.. From the point of view of the photosphere evolution, the two very smooth spectra, which are continuum blue flux dominated do not show any strong absorption features, unlike those seen in the spectra of SN 2021lwz after maximum light (see Figure 3). This observational fact, combined with the analysis of their light curves using the magnetar model (for AT 2020xnd see Table 2 in Omand & Sarin 2024), suggests that AT 2020xnd and SN 2021lwz differ in nature, and do not pertain to the same class of transients.
@@ 5.3 USSN
SN 2014ft has been observed, studied and classified by De et al. (2018) as a USSN that likely formed a compact neutron star binary. It is recognised as the first ever detection of a USSN (Hijikawa et al. 2019), while previous studies of fast evolving Type Ic-BL SNe like SN 2005ek (Drout et al. 2013) and SN 2010X (Kasliwal et al. 2010) pointed-out the existence of this class of objects. Posterior to these works, SN 2019dge (or ZTF18abfcmjw) was added to this class of transients as discussed by Yao et al. (2020).
@@
SN 2014ft is modeled using the Redback (Sarin et al. 2024) magnetar-driven supernova model in Omand & Sarin (2024). SN 2021lwz and SN 2014ft have very similar light curves both peaking at absolute magnitudes close to -20.1 mag (Figure 16), and similar fast rising times of order 6–7 days (see Table 3. Assuming the magnetar-driven supernova model is appropriate to describe both explosions, the posterior parameters from the light curve modeling both infer very low ejecta of the same order ( ), and supernovae explosion energies of the same order ( erg ). The main difference is that SN 2021lwz has an initial spin-down luminosity about 100 times higher than the one of SN 2014ft, and a spin-down timescale about 10 times faster. The braking index of both explosions being of the same order (), this results in SN 2021lwz having a total rotational energy an order of magnitude higher than SN 2014ft.
@@
During the first 30 hours after explosion SN 2014ft showed He ii 4686 Å lines in 3 spectra (see Figure3 A in De et al. 2018). Such early spectra are not available for SN 2021lwz therefore we don’t know if a similar feature was present very early in its photosphere. A similar feature is not detected in the earliest spectra available on SN 2021lwz obtained about 5 to 7 days after the explosion (see Figure 3). Hints of He i, but not of He ii, are only suggested by the modeling analysis in the post peak spectra of SN 2021lwz. At maximum light the spectra of SN 2021lwz and SN 2014ft show common ionization atomic lines from Ca ii, Mg ii, Fe ii, Ti ii, Si ii, C ii, and O i (see our SYN++ analysis and Figure 5B in De et al. 2018) but absorption lines of P iii (while not unambiguous) and Na i observed in the spectrum of SN 2021lwz are not detected in the spectrum of SN 2014ft. The distinct temperatures of the photosphere estimated for each supernova close to maximum light, with K for SN 2021lwz (this work), and K for SN 2014ft (Figure 4B in De et al. 2018), may explain this difference. Before and near maximum light SN 2021lwz has a steeper continuum than SN 2014ft, showing that it is hotter and likely coming from a more energetic event, an inference consistent with the one provided by the modeling of the light curves. In the near-maximum spectra, around 5000 Å, SN 2014ft shows a ‘basin shaped’ feature similarly to SN 2015bn, while this is missing from the spectra of SN 2021lwz, here again most probably because of the difference in the temperatures. The post-maximum spectral lines are somewhat more similar, but SN 2014ft is still cooler than SN 2021lwz.
@@
All the arguments put together it is not clear if SN 2021lwz has been ultra-stripped or not, the main argument in favor of an ultra-stripped process being that it has a low ejecta mass, as does have SN 2014ft.
@@ 5.4 Typical H-poor stripped envelope SNe
Hydrogen-poor, stripped envelope, core-collapse supernovae (SESNe), of Types Ib, Ic and Ic-BL, are well studied in the optical (e.g. Drout et al. 2011; Taddia et al. 2019) and the NIR (e.g. Shahbandeh et al. 2022). As has been discussed previously our post-peak optical spectral modeling hints at a marginal detection of He i, and the possible 1.0830 m line. Possible connections and a continuity between typical Type Ib and Ic has been discussed with no clear results about residual detections of He i in optical spectra (e.g. Hachinger et al. 2012; Piro & Morozova 2014; Prentice et al. 2018; Dessart et al. 2020; Williamson et al. 2021). More recently, Shahbandeh et al. (2022) presented the analysis of 75 NIR spectra of 34 SESNe with the aim to discuss NIR strong He i lines: m and m. In all spectra, except a few outliers, most of the absorption lines features observed close to 1 m can be explained by a combination of strong lines (C I m, He i m, and Mg II m), and the authors concluded that there is no distinct difference between He-poor and He-rich groups in this wavelength range. On the other hand, by analysing the 2 m features the authors showed one can clearly discriminate He-rich (optical IIb/Ib) from He-poor (optical Ic/Ic-BL) SESNe and computed SESNe templates using a Principal Component Analysis. The +13 days templates of such He-rich and He-poor SNe are shown in Figure 4 for illustration. A comparison with the +12.6 days NIR spectrum of SN 2021lwz, which optical spectral modeling points out to a Type Ic-like, disfavours the presence of residual He i. Actually, in the vicinity of the He i 2 m feature, the spectrum of SN 2021lwz looks flatter than the He-poor template, as looks the spectrum of the peculiar Type Ic-BL, SN 1998bw.
Additional arguments making SN 2021lwz peculiar with respect to SESNe are the following. It is more luminous at peak than the majority of these objects (see for example our plot in Figure 1 for a direct comparisons with Type Ic and Ic-BL absolute magnitudes). It has a low mass ejecta (as inferred from our multi-band modeling), , which is in contradiction with typical 56Ni ejectae masses of M and 0.5 for SNe Ib/c and SN Ic-BL, respectively (Drout et al. 2011), as well as with the ejectae masses expected for these two Types of SNe, of order M and 5 , respectively (Drout et al. 2011; Taddia et al. 2019).
5.5 Type Ib SLSN
@@ 5.6 Comparisons with peculiar Type Ic-BL SN 1998bw
Before and close to maximum light the photosphere of SN 2021lwz looks more similar to a normal Type Ic supernova than to any other Type, with C ii, O i, Na i, Mg ii, Si ii, P iii, Ca ii, Ti ii and Fe ii lines identified in its earliest spectrum (see Section 4.1), and the velocity of the spectral features and of the photosphere are close to those of a Type Ic. Broader optical spectral features observed after peak suggest that SN 2021lwz could have evolved to a Type Ic-BL. Actually, before maximum light, the SYN++ modeling shows that the photosphere velocity of SN 2021lwz is not extraordinarily high ( km/s), and the broadness parameter (see values displayed in Table 9) are pointing on a normal Type Ic, to the limit of a Ic-BL. However, after maximum light, the input parameter value had to be higher to adjust the formation of several broader spectral features, making SN 2021lwz more similar to a Type Ic-BL SN. Some comparisons of post-peak spectra of SN 2021lwz and peculiar Type Ic-BL SN 1998bw are shown in Figure 18, attesting the likely BL nature of SN 2021lwz in these phases, but the best match is obtained with LSQ12dlf which is a SLSN-I showing type Ic BL after peak. The match with the +14 days and +16 days spectra of SN 1998bw are relatively good above but show some differences at lower wavelengths. Some attempts to fit the -3.4 days Keck spectrum of SN 2021lwz with early spectra of SN 1998bw proved unsuccesful. Some differences between the NIR spectra of both transients can also seen in Figure 4.
In conclusion, while SN 2021lwz and SN 1998bw both appear to be peculiar Type Ic/Ic-BL transients, their discrepancies suggest they are two different types of objects.
@@ 5.7 Comparisons with fast evolving luminous Type Ic-BL iPTF16asu
iPTF 16asu whose origin is not yet well understood has been first discussed by Whitesides et al. (2017). It is a luminous, rapidly evolving, stripped-envelope SN peaking at mag. With a rest-frame rise time estimated as 4 days it appears as a faster evolving event than SN 2021lwz. Contrary to SN 2021lwz which looks like a Type Ic SN before peak, iPTF 16asu shows featureless blue continuum, but after maximum light SN 2021lwz and iPTF 16asu both show Ic-BL features. In Appendix H we present the fit of the multiband light curve of iPTF 16asu using the same software, model, and priors as for SN 2021lwz (see Section 4.3 for details). The light curve fit is shown in Figure 20 and posterior is shown in Figure 21. From our modeling, the estimated SN explosion energy of iPTF 16asu is about 30 times stronger than the one of SN 2021lwz, which could explain the optical spectral differences before peak.
@@
More in depth light curve modeling of iPTF 16asu has been discussed by Wang et al. (2019). In Section 4.2 we discuss the difficulty to model the light curve of SN 2021lwz using the Arnett model, similarly the analysis proposed by these authors underlined the difficulty to model the light curve of iPTF 16asu with radioactive heating (56Ni-only), and also the magnetar 56Ni models. Our modeling of iPTF 16asu and SN 2021lwz includes a fraction of Nickel in the ejecta mass (see Equation 9 in Omand & Sarin 2024), which given the total ejecta mass estimate and , of each object, respectively, means the Nickel mass would be excessively low. Energetically speaking, we find that these low ejecta masses are both associated to rotational energies of the same order (). The light curves of both events are very similar and peak at absolute magnitudes close to -20.3 mag. They also have fast rising times sharing a common parameter space ( days; see Table 3 and Figure 1). The g-band data plotted in Figure 16 show the steep increase of iPTF 16asu, while the r-band data show a decreasing ligth curve very similar to the one of SN 2021lwz.
@@
The interpretation of the results obtained by Wang et al. (2019), and additional modeling of the light curves has been proposed by Wang & Gan (2022). The construction of the theoretical bolometric light curve is revised and the authors find that the magnetar plus 56Ni model can now account for the photometry of iPTF 16asu with a peak luminosity inferred from the theoretical light curve of erg. The best fit value for the ejecta mass is and the best fit parameter of the Nickel mass is . Some results very close to the best fit parameters obtained with our modeling of the multi-light curve of SN 2021lwz. All these results possibly make SN 2021lwz part of a class of rare rapidly evolving ( week), luminous type Ic-BL SNe ( mag or slightly higher), like iPTF 16asu (see Wang & Gan 2022), even though, SN 2021lwz is an intermediate velocity ( km s-1) type Ic/Ic-BL object, while iPTF 16asu is a high velocity type Ic-BL after peak.

### v2

@@ 5 Discussion
Based on our previous modelling and analysis it is not clear what is the nature of SN 2021lwz. This very fast evolving explosion (t days, rest-frame), with a low ejecta mass ( ), is drawn from a very low-brightness ( mag), low mass () dwarf galaxy of SFR () about ten times larger than the typical star-forming galaxies.
| Object | R.A. | Decl. | Type | z | trise,rf (a) | band | (b) | (c) | References |
@@
To discuss SN 2021lwz in a broader context, in Figure 15 we show the rest-frame light curves of the different types of transients listed in Table 3; namely SN 2018bgv (SLSN-I), SN 1998bw (peculiar Ic-BL), AT 2020xnd (LFBOT), SN 2014ft (USSN), iPTF16asu (SL Ic-BL) and SN 2018gep (SL Ic-BL). As it is done for SN 2021lwz, if the photometry is not already corrected for the extinction contribution from our Galaxy, it is done using Schlafly and Finkbeiner (2011). As much as possible the filters were chosen by taking into account the redshift of each source so that the rest-frame wavelength ranges covered overlap strongly with each other.
@@ 5.1 H-poor SLSNe
SN 2018bgv is the fastest evolving ( days rest-frame) Type-W SLSN-I reported in the literature. It peaks close to mag. It is therefore very interesting for comparisons with SN 2021lwz. While we didn’t model it with our magnetar model, the light curve and a spectral sequence have been studied in detail by Lunnan et al. (2020). From black-body fits analysis of the light curve the authors report a photosphere temperature close to 11,000 K and a black-body radius close to 2 cm, a few days after peak, i.e. a photopshere with a radius about twice larger than, and a temperature about similar to, those of SN 2021lwz photosphere at similar phases. Close to peak the velocity of the photosphere of SN 2018bgv appears to be about 1.5 times the velocity estimated for SN 2021lwz. A fit of the light curve with the magnetar spin-down model using mosfit infers an ejecta of mass, , i.e. about 5 times higher than the one of SN 2021lwz as from our modelling estimate, hence suggesting the conditions behind the explosions of SN 2018bgv and SN 2021lwz are different.
As discussed in Section 3.1 SN 2018bgv provides a relatively good match with SN 2021lwz close to maximum light, but this is obtained at a lower redshift z than the redshift of SN 2021lwz. This is mainly because SNID tries to adjust the archetypal W-shape feature attributed to O ii, before and close to maximum light, observed in many Type I SLSNe. Indeed, and as discussed in Section 4.1, for H-poor SLSNe the typical absorption spectral features expected to be observed before maximum light, between 4000 and 5000 Å (see Könyves-Tóth and Seli 2023, Figure 1), in Type 15bn (e.g. Fe iii) and in Type W (e.g. O ii), are not detected in the pre-maximum spectra of SN 2021lwz. Instead, C ii, O i, Na i, Mg ii, Si ii, Ca ii, Ti ii and Fe ii are included to reproduce the shape of these spectra, and this despite the fact that from the estimates of the photosphere temperature K and the photosphere velocities kms before maximum light, one would expect features from a W-Type SLSN-I (see Könyves-Tóth 2022, Figure 7). We point out, however, that the conditions for these lines to appear are likely limited to a more stringent range of temperature K once spectral modelling effects, which are not included in the SYN++ modelling, are taken into account (see Saito et al. 2024). However, the study provided by Aamer et al. (2025), obtained on a sample larger than the one used by Könyves-Tóth and Seli (2023) suggests a continuum of spectral properties, hence that the differences of spectral profiles in type I SLSNe may not correspond to physically different type of events. In this larger sample Fe ii lines appear to be an important spectral marker formed near the photosphere as it is also found for SN 2021lw.
@@ 5.2 LFBOTs
The LFBOT AT 2020xnd and SN 2021lwz are peaking at about similar absolute magnitudes (see Figure 15), but AT 2020xnd is a faster evolving event than SN 2021lwz, mainly because it has a higher initial pulsar luminosity, assuming a magnetar model. Two spectra of AT 2020xnd at phases close to +5 days and +7 days after maximum light are publicly available in the WISeREP repository 181818URL to AT 2020xnd public spectra in WISeREP, https://www.wiserep.org/object/18199.. From the point of view of the photosphere evolution, the two very smooth spectra, which are continuum blue flux dominated do not show any strong absorption features, unlike those seen in the spectra of SN 2021lwz after maximum light (see Figure 3). This observational fact, combined with the analysis of their light curves using the magnetar model (for AT 2020xnd see Table 2 in Omand and Sarin 2024), suggests that AT 2020xnd and SN 2021lwz differ in nature, and do not pertain to the same class of transients. This conclusion is sustained by a visual comparison of some spectra of SN 2021lwz with some of SN 2020xnd and SN 2018cow as shown in Figure 16. The spectral evolution of SN 2021lwz show absorption features stronger than those observed in the spectra of AT 2020xnd and AT 2018cow which, in comparison, look almost featureless.
@@ 5.3 USSN
SN 2014ft has been observed, studied and classified by De et al. (2018) as a USSN that likely formed a compact neutron star binary. It is recognised as the first ever detection of a USSN (Hijikawa et al. 2019), while previous studies of fast evolving Type Ic-BL SNe like SN 2005ek (Drout et al. 2013) and SN 2010X (Kasliwal et al. 2010) pointed-out the existence of this class of objects. Posterior to these works, SN 2019dge (or ZTF18abfcmjw) was added to this class of transients as discussed by Yao et al. (2020). More recently SN 2023zaw has been shown to be a fast evolving USSN of very low ejecta mass ( ) (Das et al. 2024; Moore et al. 2025).
@@
SN 2014ft is modelled using the Redback (Sarin et al. 2024) magnetar-driven supernova model in Omand and Sarin (2024). SN 2021lwz and SN 2014ft have very similar light curves both peaking at absolute magnitudes close to -20.1 mag (Figure 15), and similar fast rising times of order 6–7 days (see Table 3. Assuming the magnetar-driven supernova model is appropriate to describe both explosions, the posterior parameters from the light curve modelling both infer very low ejecta of the same order ( ), and supernovae explosion energies of the same order ( erg ). The main difference is that SN 2021lwz has an initial spin-down luminosity about 100 times higher than the one of SN 2014ft, and a spin-down timescale about 10 times faster. The braking index of both explosions being of the same order (), this results in SN 2021lwz having a total rotational energy an order of magnitude higher than SN 2014ft.
@@
During the first 30 hours after explosion SN 2014ft showed He ii 4686 Å lines in 3 spectra (see Figure3 A in De et al. 2018). Such early spectra are not available for SN 2021lwz therefore we don’t know if a similar feature was present very early in its photosphere. A similar feature is not detected in the earliest spectra available on SN 2021lwz obtained about 5 to 7 days after the explosion (see Figure 3). Hints of He i, but not of He ii, are only suggested by the modelling analysis in the post peak spectra of SN 2021lwz. At maximum light the spectra of SN 2021lwz and SN 2014ft show common ionization atomic lines from Ca ii, Mg ii, Fe ii, Ti ii, Si ii, C ii, and O i (see our SYN++ analysis and Figure 5B in De et al. 2018) but absorption lines of P iii (while not unambiguous) and Na i observed in the spectrum of SN 2021lwz are not detected in the spectrum of SN 2014ft. The distinct temperatures of the photosphere estimated for each supernova close to maximum light, with K for SN 2021lwz (this work), and K for SN 2014ft (Figure 4B in De et al. 2018), may explain this difference. Before and near maximum light SN 2021lwz has a steeper continuum than SN 2014ft, showing that it is hotter and likely coming from a more energetic event, an inference consistent with the one provided by the modelling of the light curves. In the near-maximum spectra, around 5000 Å, SN 2014ft shows a ‘basin shaped’ feature similarly to SN 2015bn, while this is missing from the spectra of SN 2021lwz, here again most probably because of the difference in the temperatures. The post-maximum spectral lines are somewhat more similar, but SN 2014ft is still cooler than SN 2021lwz.
@@
All the arguments put together it is not clear if SN 2021lwz has been ultra-stripped or not, the main argument in favour of an ultra-stripped process being that it has a low ejecta mass, as does have SN 2014ft.
@@ 5.4 Typical H-poor stripped envelope SNe
Hydrogen-poor, stripped envelope, core-collapse supernovae (SESNe), of Types Ib, Ic and Ic-BL, are well studied in the optical (e.g. Drout et al. 2011; Taddia et al. 2019) the NIR (e.g. Shahbandeh et al. 2022), and at radio wavelengths (Srinivasaragavan et al. 2024). As has been discussed previously our post-peak optical spectral modelling hints at a marginal detection of He i, and the possible 1.0830 m line. Possible connections and a continuity between typical Type Ib and Ic has been discussed with no clear results about residual detections of He i in optical spectra (e.g. Hachinger et al. 2012; Piro and Morozova 2014; Prentice et al. 2018; Dessart et al. 2020; Williamson et al. 2021). More recently, Shahbandeh et al. (2022) presented the analysis of 75 NIR spectra of 34 SESNe with the aim to discuss NIR strong He i lines: m and m. In all spectra, except a few outliers, most of the absorption lines features observed close to 1 m can be explained by a combination of strong lines (C I m, He i m, and Mg II m), and the authors concluded that there is no distinct difference between He-poor and He-rich groups in this wavelength range. On the other hand, by analysing the 2 m features the authors showed one can clearly discriminate He-rich (optical IIb/Ib) from He-poor (optical Ic/Ic-BL) SESNe and computed SESNe templates using a Principal Component Analysis. The +13 days templates of such He-rich and He-poor SNe are shown in Figure 4 for illustration. A comparison with the +12.6 days NIR spectrum of SN 2021lwz, which optical spectral modelling points out to a Type Ic-like, disfavours the presence of residual He i. Actually, in the vicinity of the He i 2 m feature, the spectrum of SN 2021lwz looks flatter than the He-poor template, as looks the spectrum of the peculiar Type Ic-BL, SN 1998bw.
Additional arguments making SN 2021lwz peculiar with respect to SESNe are the following. It is more luminous at peak than the majority of these objects (see for example our plot in Figure 1 for a direct comparisons with Type Ic and Ic-BL absolute magnitudes). It has a low mass ejecta (as inferred from our multi-band modelling), , which is in contradiction with typical 56Ni ejecta masses of M and 0.5 for SNe Ib/c and SN Ic-BL, respectively (Drout et al. 2011), as well as with the ejecta masses expected for these two Types of SNe, of order M and 5 , respectively (Drout et al. 2011; Taddia et al. 2019). More recently Rodríguez et al. (2023) studied 191 stripped-envelope SNe, including 50 SNe Ib and 96 SNe Ic, of which 26 are SNe Ic-BL. Compared to that SN Ib/Ic/Ic-BL sample, SN 2021lwz is at least ten times more luminous at peak than the average (2.7 and 4.5 erg s-1 for Ib and Ic/Ic-BL, respectively). Its ejecta mass is about ten times lower than the average (2.6 ), and also lower than the minimum ejecta mass (0.5 ), making it a low occurrence event in the bottom tail of the distribution.
@@ 5.5 Comparisons with peculiar Type Ic-BL SN 1998bw
Before and close to maximum light the photosphere of SN 2021lwz looks more similar to a normal Type Ic supernova than to any other Type, with C ii, O i, Na i, Mg ii, Si ii, Ca ii, Ti ii and Fe ii lines identified in its earliest spectrum (see Section 4.1), and the velocity of the spectral features and of the photosphere are close to those of a Type Ic. Slightly broader optical spectral features observed after peak suggest that SN 2021lwz could have evolved closer to a Type Ic-BL. Actually, before maximum light, the SYN++ modelling shows that the photosphere velocity of SN 2021lwz is not extraordinarily high ( km/s), and the broadness parameter (see values displayed in Table 9) are pointing on a normal Type Ic. After maximum light, the input parameter value had to be higher to adjust the formation of several broader spectral features, making SN 2021lwz may be closer to the limit of a Ic-BL. But still, after peak, the best match is obtained with LSQ12dlf which is a SLSN-I showing velocity lines below the 30000 km/s seen in classic SNe Ic-BL. The match with the +14 days and +16 days spectra of SN 1998bw are not too bad above but are relatively poor at lower wavelengths. Some attempts to fit the -3.4 days Keck spectrum of SN 2021lwz with early spectra of SN 1998bw proved unsuccessful. Some differences between the NIR spectra of both transients can also seen in Figure 4. All in all, SN 2021lwz looks more like a SLSN showing the Ic-SL Type after peak.
@@ 5.6 Comparisons with fast evolving luminous Type Ic-BL iPTF16asu
iPTF 16asu whose origin is not yet well understood has been first discussed by Whitesides et al. (2017). It is a luminous, rapidly evolving, stripped-envelope SN peaking at mag. With a rest-frame rise time estimated as 4 days it appears as a faster evolving event than SN 2021lwz. Contrary to SN 2021lwz which looks like a Type Ic SN before peak, iPTF 16asu shows a relatively featureless blue continuum, but contrary to SN 2021lwz, iPTF 16asu shows Ic-BL features after peak (see some spectra shown in Figure 17 for comparison).
In Appendix I we present the fit of the multi-band light curve of iPTF 16asu using the same software, model, and priors as for SN 2021lwz (see Section 4.3 for details). The light curve fit is shown in Figure 22 and posterior is shown in Figure 23. From our modelling, the estimated SN explosion energy of iPTF 16asu is about 30 times stronger than the one of SN 2021lwz, which could explain the optical spectral differences before peak.
@@
More in depth light curve modelling of iPTF 16asu has been discussed by Wang et al. (2019). In Section 4.2 we discuss the difficulty to model the light curve of SN 2021lwz using the Arnett model, similarly the analysis proposed by these authors underlined the difficulty to model the light curve of iPTF 16asu with radioactive heating (56Ni-only), and also the magnetar 56Ni models. Our modelling of iPTF 16asu and SN 2021lwz includes a fraction of Nickel in the ejecta mass (see Equation 9 in Omand and Sarin 2024), which given the total ejecta mass estimate and , of each object, respectively, means the Nickel mass would be excessively low. Energetically speaking, we find that these low ejecta masses are both associated to rotational energies of the same order (). The light curves of both events are very similar and peak at absolute magnitudes close to -20.3 mag. They also have fast rising times sharing a common parameter space ( days; see Table 3 and Figure 1). The g-band data plotted in Figure 15 show the steep increase of iPTF 16asu, while the r-band data show a decreasing light curve very similar to the one of SN 2021lwz.
@@
The interpretation of the results obtained by Wang et al. (2019), and additional modelling of the light curves has been proposed by Wang and Gan (2022). The construction of the theoretical bolometric light curve is revised and the authors find that the magnetar plus 56Ni model can now account for the photometry of iPTF 16asu with a peak luminosity inferred from the theoretical light curve of erg. The best fit value for the ejecta mass is and the best fit parameter of the Nickel mass is . Some results very close to the best fit parameters obtained with our modelling of the multi-light curve of SN 2021lwz. All these results possibly make SN 2021lwz part of a class of rare rapidly evolving ( week), luminous type Ic-BL SNe ( mag or slightly higher), like iPTF 16asu (see Wang and Gan 2022), even though, SN 2021lwz is an intermediate velocity ( km s-1) type Ic/Ic-BL object, while iPTF 16asu is a high velocity type Ic-BL after peak.